# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part II: posterior mean approximation

### v2

@@ B.2 Proofs of Section 5
We show the forward inclusions next. Suppose that $A\in\mathscr{M}^{(i)}_{r}$ for $i=1$ or $i=2$. By Theorem 3.2(ii), $AY-m_{\textup{pos}}(Y)\in\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$ with probability 1. Since $m_{\textup{pos}}(Y)\in\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$ with probability 1 by (3a), this implies $AY\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1. Now fix $i=2$. By Lemma A.15, $AY$ is a Gaussian measure with covariance $A\mathcal{C}_{\textup{y}}A^{*}$, where $\mathcal{C}_{\textup{y}}$ is the covariance of $Y$. By [3, Theorem 2.4.7] or [17, Proposition 4.45], the Cameron–Martin space of a Gaussian measure is contained in every measurable linear subspace of full measure. Thus, since $AY\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1, the Cameron–Martin space of $AY$, which is $\operatorname{ran}{(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}}$, is contained in $\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Because $A$ has finite rank, $A\mathcal{C}_{\textup{y}}A^{*}$ has finite rank and therefore $\operatorname{ran}{A\mathcal{C}_{\textup{y}}A^{*}}=\operatorname{ran}{(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}}$, by Lemma A.6 applied to $A\leftarrow(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}$. Furthermore, by Lemma A.6 applied to $A\leftarrow A\mathcal{C}_{\textup{y}}^{1/2}$ and invertibility of $\mathcal{C}_{\textup{y}}$, we have $\operatorname{ran}{A}=\operatorname{ran}{A\mathcal{C}_{\textup{y}}}^{1/2}=\operatorname{ran}{A\mathcal{C}_{\textup{y}}^{1/2}(A\mathcal{C}_{\textup{y}}^{1/2})^{*}}=\operatorname{ran}{A\mathcal{C}_{\textup{y}}A^{*}}$. As a consequence, $\operatorname{ran}{A}=\operatorname{ran}{(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. This shows the forward inclusion for $i=2$. Finally, let $i=1$. Thus, $A=(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ for some $B\in\mathcal{B}_{00,r}(\mathcal{H})$. Since we just showed that $AY\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1, and since $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$, it follows that $BG^{*}\mathcal{C}_{\textup{obs}}^{-1}Y\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1. By replacing $A$ with $BG^{*}\mathcal{C}_{\textup{obs}}^{-1}$ in the argument for the case where $i=2$, we obtain $\operatorname{ran}{BG^{*}}\mathcal{C}_{\textup{obs}}^{-1}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Since $\operatorname{ran}{G^{*}}$ is finite-dimensional, it is closed. Using that $\mathcal{C}_{\textup{obs}}$ is invertible, this implies $B(\ker{G}^{\perp})=B(\operatorname{ran}{G^{*}})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ by Lemma A.4. This shows the forward inclusion for $i=1$.
(ii) By Lemma 5.3(i), $S_{\textup{pos}}$ is injective and $\operatorname{ran}{S_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$. Thus, $\operatorname{rank}\left({S_{\textup{pos}}\widetilde{A}}\right)=\operatorname{rank}\left({\widetilde{A}}\right)$ and $\operatorname{ran}{S_{\textup{pos}}\widetilde{A}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}}$ for every $\widetilde{A}\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H})$. By (24b), $S_{\textup{pos}}\widetilde{\mathscr{M}}^{(2)}_{r}=\{A\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}):\ \operatorname{ran}{A}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\}=\mathscr{M}^{(2)}_{r}$. This shows the result for $i=2$. For $i=1$, first let $A\in\mathscr{M}^{(1)}_{r}$. By (24a), this implies $A=(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ for some $B\in\mathcal{B}_{00,r}(\mathcal{H})$ with $B(\ker{G}^{\perp})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Let $\widetilde{B}\coloneqq S_{\textup{pos}}^{-1}BP_{\ker{G}^{\perp}}$, where $P_{\ker{G}^{\perp}}$ denotes the orthogonal projector onto $\ker{G}^{\perp}$. Then $\widetilde{B}$ is well-defined, because $\operatorname{ran}{BP_{\ker{G}^{\perp}}}=B(\ker{G}^{\perp})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{dom}{S_{\textup{pos}}^{-1}}$ by Lemma 5.3(i). Furthermore, $\operatorname{rank}\left({\widetilde{B}}\right)\leq\operatorname{rank}\left({B}\right)\leq r$ and $S_{\textup{pos}}\widetilde{B}=BP_{\ker{G}^{\perp}}$. Hence $S_{\textup{pos}}\widetilde{B}G^{*}=BG^{*}$ by Lemma A.4, showing $A=S_{\textup{pos}}(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B})G^{*}\mathcal{C}_{\textup{obs}}^{-1}$. Thus, $A\in S_{\textup{pos}}\widetilde{\mathscr{M}}^{(1)}_{r}$. For the reverse inclusion, let $A\in S_{\textup{pos}}\widetilde{\mathscr{M}}^{(1)}_{r}$. That is, let $A=S_{\textup{pos}}(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B})G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ for some $\widetilde{B}\in\mathcal{B}_{00,r}(\mathcal{H})$. Then $A=(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}$, where $B\coloneqq S_{\textup{pos}}\widetilde{B}$ satisfies $\operatorname{rank}\left({B}\right)=\operatorname{rank}\left({\widetilde{B}}\right)\leq r$ and $B(\ker{G}^{\perp})\subset\operatorname{ran}{B}\subset\operatorname{ran}{S_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. By (24a), this shows that $A\in\mathscr{M}^{(1)}_{r}$.
@@
and then set $A_{r}^{\textup{opt},(2)}\coloneqq S_{\textup{pos}}\widetilde{A}_{r}^{\textup{opt},(2)}$ using Proposition 5.5(iii). Note that $I^{\dagger}=I$, that $S_{y}^{\dagger}=S_{\textup{y}}^{-1}$ by Lemma 5.3(i), and that $(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2})_{r}\coloneqq\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$ is a rank-$r$ truncated SVD of $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ by (20). Since $I\in\mathcal{B}(\mathcal{H})$ and $S_{\textup{y}}\in\mathcal{B}(\mathbb{R}^{n})$ have closed range, and since $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ has finite rank and is thus Hilbert–Schmidt, we may apply Theorem 5.7 with $\mathcal{H}_{i}\leftarrow\mathbb{R}^{n}$ for $i\in\{1,2\}$, $\mathcal{H}_{i}\leftarrow\mathcal{H}$ for $i\in\{3,4\}$, $T\leftarrow I$, $S\leftarrow S_{\textup{y}}$, $M\leftarrow\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ to find
@@
where $w_{i}$ is the eigenvector corresponding to the eigenvalue $\lambda_{i}$ given by Proposition 3.4 and $\varphi_{i}$ is the right singular vector corresponding to $\lambda_{i}$ in (20). By [13, Theorem 2.8], the Moore–Penrose inverse of $\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$ is given by $\sum_{i=1}^{n}\sqrt{\frac{1+\lambda_{i}}{-\lambda_{i}}}\varphi_{i}\otimes w_{i}$. Furthermore, the Moore–Penrose inverse of a composition of bounded operators is the composition in reverse order of the Moore–Penrose inverses of these operators, e.g. [18, eq. (3.23)]. Since $\mathcal{C}_{\textup{pr}}^{-1/2}$ and $I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}$ are boundedly invertible by Lemma A.7, it thus holds that the bounded operator
@@
has Moore–Penrose inverse equal to $S$. Because [2, Theorem 9.2(f)] implies that $(\mathfrak{S}^{\dagger})^{\dagger}=\mathfrak{S}$ for any bounded operator $\mathfrak{S}$, the operator in the display above is equal to $S^{\dagger}$. Furthermore, by [13, eq. (2.12)], $P_{\ker{S}^{\perp}}=S^{\dagger}S$, showing that $P_{\ker{S}^{\perp}}=\sum_{i=1}^{n}\varphi_{i}\otimes\varphi_{i}.$ Next, we compute for the given choice of $M$,
@@ B.3 Proofs of Section 7
Recall that ${H}$ defined in (2) is the Hessian of the negative log-likelihood of (1). Analogously, let $\widetilde{H}$ denote the Hessian of the negative log-likelihood of (29) with the given $P$. That is, upon replacement of $G$ with $GP$ in (2), we obtain $\widetilde{H}$. Hence, orthonormality of $(\varphi_{i})_{i}$ implies
@@ Appendix C Examples
Let us now assume that $\kappa$ is bounded and symmetric, and satisfies $\int\kappa(s,t)h(s)h(t)\geq 0$ for all $h\in\mathcal{H}$. Hence, $T_{\kappa}$ is self-adjoint and nonnegative. Then by Mercer’s theorem, [19, Theorem 3.a.1], we have $\kappa(s,t)=\sum_{i=1}^{\infty}b_{i}f_{i}(s)f_{i}(t)$, where the series converges absolutely and uniformly for almost every $(t,s)$. Here, $(b_{i})_{i}$ is a nonnegative sequence converging to zero and $(f_{i})_{i}$ is an ONB of $\mathcal{H}$ consisting of bounded functions. Furthermore, we may write $T_{\kappa}=\sum_{i}b_{i}f_{i}\otimes f_{i}$. For simplicity, we assume that the eigenvectors $(e_{i})_{i}$ of the prior covariance and the eigenfunctions $(f_{i})_{i}$ of the kernel are the same. One can verify that, with $a_{k,j}\coloneqq\langle f_{k},1_{[t_{j},t_{j+1}]}\gamma\rangle$, we have $\langle T_{\kappa}h,1_{[t_{i},t_{i+1}]}\gamma\rangle=\sum_{j}b_{j}a_{j,i}\langle f_{j},h\rangle$ and $\int\kappa(t,\cdot)1_{[t_{i},t_{i+1}]}(t)\gamma(t)\operatorname{d}\!{t}=\sum_{k}b_{k}a_{k,i}f_{k}$. Thus, $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}z=\sum_{i=1}^{n}\sum_{j}z_{i}b_{j}c_{j}a_{j,i}f_{j}$ for $z\in\mathbb{R}^{n}$. Furthermore, $G^{*}G=\sum_{i=1}^{n}\sum_{j,k}b_{j}b_{k}a_{j,i}a_{k,i}f_{k}\otimes f_{j}$ and hence the prior-preconditioned Hessian now takes the form
@@
$\partial_{t}u-\partial_{xx}u$ $=0,$ in ($\text{in }(0,1)\times(0,T),$
[…]
$u(0,\cdot)=u(1,\cdot)$ $=0,$ on ($\text{on }(0,T],$
where the true initial condition $x^{\dagger}$ is unknown and where we impose a homogenous Dirichlet spatial boundary condition. We assume that the data consists of a noisy observation of $u$ at the observation coordinates $(x_{i},t_{i})_{i=1}^{n}\subset(0,1)\times(0,T]$, where we assume i.i.d. standard Gaussian noise. The aim is to reconstruct the initial condition $x^{\dagger}$ from the data $y$. This problem is similar to [34, Example 3.5] and [15, Section 4.2], but in this example we do not observe the temperature field over the entire spatial domain at finitely many times. Instead, we observe the temperature only at finitely many space-time points $(x_{i},t_{i})_{i=1}^{n}$. Furthermore, [15, Section 4.2] considers periodic boundary conditions instead of Dirichlet boundary conditions. We take the Bayesian perspective by considering $x^{\dagger}$ as an $\mathcal{H}$-valued random variable $X$ with centered Gaussian distribution $\mu_{\textup{pr}}$. Below, we choose an explicit form of the prior covariance $\mathcal{C}_{\textup{pr}}$ as a negative power of the Laplacian.
To write this problem in the formulation of Section 2, we define $\mathcal{H}\coloneqq L^{2}((0,1))$. Let us denote by $H^{1}((0,1))$ the Sobolev space of square-integrable functions $h$ on $(0,1)$ that have a square-integrable weak derivative $\partial_{x}h$, which is a Hilbert space with the inner product $\langle h_{1},h_{2}\rangle_{1}\coloneqq\langle h_{1},h_{2}\rangle+\langle\partial_{x}h_{1},\partial_{x}h_{2}\rangle$, $h_{1},h_{2}\in H^{1}((0,1))$. By [14, Theorem 5.6.5], we have the continuous embedding $H^{1}((0,1))\subset C([0,1])$, where $C([0,1])$ denotes the space of continuous functions on $[0,1]$ with the supremum norm. Hence, for any $h\in H^{1}((0,1))$ and $x\in[0,1]$, we have $\lvert h(x)\rvert\leq\lVert h\rVert_{C([0,1])}\leq c\lVert h\rVert_{1}$ for some $c>0$, so that pointwise evaluation is well-defined, linear and continuous on $H^{1}((0,1))$. Thus, $H^{1}((0,1))$ is a reproducing kernel Hilbert space. We denote the Riesz representatives of the pointwise evaluation functionals, or ‘features’, by $\{\phi(x)\in H^{1}((0,1)),\ x\in[0,1]\}$. Hence, $h(x)=\langle h,\phi(x)\rangle_{1}$ for all $x\in[0,1]$ and $h\in H^{1}((0,1))$. For our choice of spatial domain $(0,1)$, we have the following explicit form for the features, by [35, Corollary 2]:
@@
We use certain properties of $\Delta\coloneqq\partial_{xx}$, the one-dimensional Laplacian. We describe these briefly, and refer to [20, Section 5.3] for a comprehensive treatment of these properties and their relation to the heat equation. By [4, Theorem 8.22], we can write $\Delta h=-\sum_{i}a_{i}\langle h,e_{i}\rangle e_{i}$ for $h\in\operatorname{dom}{\Delta}=\{h\in L_{2}((0,1)):\ \sum_{i}a_{i}^{2}\langle h,e_{i}\rangle^{2}<\infty\}$, where $\lim_{i}a_{i}=\infty$ and $(e_{i})_{i}$ is an ONB on $\mathcal{H}$. In fact, by the example on [4, p. 232], we have $a_{i}=i^{2}\pi^{2}$ and $e_{i}(x)=\sqrt{2}\sin(i\pi x)$ for our choices of spatial domain $(0,1)$ and boundary conditions. Now, one can define the self-adjoint operator $\exp(t\Delta)\in\mathcal{B}_{0}(\mathcal{H})$ by $\exp(t\Delta)=\sum_{i}\exp(-ta_{i})e_{i}\otimes e_{i}$. It holds that $\ker{\exp(t\Delta)}=\{0\}$ and $\operatorname{ran}{\exp(t\Delta)=\mathcal{H}}$. The diagonalisation of the Laplacian is compatible with $H^{1}_{0}((0,1))$ in the sense that $H^{1}_{0}((0,1))=\{h\in\mathcal{H}:\ \sum_{i}a_{i}\langle h,e_{i}\rangle^{2}<\infty\}$ and $\langle h,k\rangle_{1}=\sum_{i}(1+a_{i})\langle h,e_{i}\rangle\langle e_{i},k\rangle$ for $h,k\in H^{1}_{0}((0,1))$. Since for any $t\in(0,T]$ and $h\in\mathcal{H}$, we have $\langle\exp(t\Delta)h,e_{i}\rangle=\exp(-a_{i}t)\langle h,e_{i}\rangle$, it follows that $\sum_{i}a_{i}\langle\exp(t\Delta)h,e_{i}\rangle^{2}\leq C(t)\sum_{i}\langle h,e_{i}\rangle^{2}$ for some $C(t)>0$, so that $\exp(t\Delta)h\in\mathcal{H}^{1}_{0}((0,1))$. Therefore, the map $h\mapsto\exp(\Delta t)h$, $\mathcal{H}\rightarrow H^{1}_{0}((0,1))$ is linear and continuous for $t\in(0,T]$. Furthermore, $\exp(\Delta t)\in\mathcal{B}(H^{1}_{0}((0,1)))$ for each $t\in[0,T]$, and $\exp(\Delta t)$ is a self-adjoint element of $\mathcal{B}(H^{1}_{0}((0,1)))$, because

### v3

@@ B.2 Proofs of Section 5
We show the forward inclusions next. Suppose that $A\in\mathscr{M}^{(i)}_{r}$ for $i=1$ or $i=2$. By Theorem 3.2(ii), $AY-m_{\textup{pos}}(Y)\in\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$ with probability 1. Since $m_{\textup{pos}}(Y)\in\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$ with probability 1 by (3a), this implies $AY\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1. Now fix $i=2$. By Lemma A.15, $AY$ is a Gaussian measure with covariance $A\mathcal{C}_{\textup{y}}A^{*}$, where $\mathcal{C}_{\textup{y}}$ is the covariance of $Y$. By [4, Theorem 2.4.7] or [22, Proposition 4.45], the Cameron–Martin space of a Gaussian measure is contained in every measurable linear subspace of full measure. Thus, since $AY\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1, the Cameron–Martin space of $AY$, which is $\operatorname{ran}{(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}}$, is contained in $\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Because $A$ has finite rank, $A\mathcal{C}_{\textup{y}}A^{*}$ has finite rank and therefore $\operatorname{ran}{A\mathcal{C}_{\textup{y}}A^{*}}=\operatorname{ran}{(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}}$, by Lemma A.6 applied to $A\leftarrow(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}$. Furthermore, by Lemma A.6 applied to $A\leftarrow A\mathcal{C}_{\textup{y}}^{1/2}$ and invertibility of $\mathcal{C}_{\textup{y}}$, we have $\operatorname{ran}{A}=\operatorname{ran}{A\mathcal{C}_{\textup{y}}}^{1/2}=\operatorname{ran}{A\mathcal{C}_{\textup{y}}^{1/2}(A\mathcal{C}_{\textup{y}}^{1/2})^{*}}=\operatorname{ran}{A\mathcal{C}_{\textup{y}}A^{*}}$. As a consequence, $\operatorname{ran}{A}=\operatorname{ran}{(A\mathcal{C}_{\textup{y}}A^{*})^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. This shows the forward inclusion for $i=2$. Finally, let $i=1$. Thus, $A=(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ for some $B\in\mathcal{B}_{00,r}(\mathcal{H})$. Since we just showed that $AY\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1, and since $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$, it follows that $BG^{*}\mathcal{C}_{\textup{obs}}^{-1}Y\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ with probability 1. By replacing $A$ with $BG^{*}\mathcal{C}_{\textup{obs}}^{-1}$ in the argument for the case where $i=2$, we obtain $\operatorname{ran}{BG^{*}}\mathcal{C}_{\textup{obs}}^{-1}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Since $\operatorname{ran}{G^{*}}$ is finite-dimensional, it is closed. Using that $\mathcal{C}_{\textup{obs}}$ is invertible, this implies $B(\ker{G}^{\perp})=B(\operatorname{ran}{G^{*}})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ by Lemma A.4. This shows the forward inclusion for $i=1$.
[…]
@@
and then set $A_{r}^{\textup{opt},(2)}\coloneqq S_{\textup{pos}}\widetilde{A}_{r}^{\textup{opt},(2)}$ using Proposition 5.5(iii). Note that $I^{\dagger}=I$, that $S_{y}^{\dagger}=S_{\textup{y}}^{-1}$ by Lemma 5.3(i), and that $(\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2})_{r}\coloneqq\sum_{i=1}^{r}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$ is a rank-$r$ truncated SVD of $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ by (20). Since $I\in\mathcal{B}(\mathcal{H})$ and $S_{\textup{y}}\in\mathcal{B}(\mathbb{R}^{n})$ have closed range, and since $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ has finite rank and is thus Hilbert–Schmidt, we may apply Theorem 5.7 with $\mathcal{H}_{i}\leftarrow\mathbb{R}^{n}$ for $\in\{1,2\}, i\mathcal{H}_{i}\leftarrow\mathcal{H}$ for $i\in\{3,4\}$, $T\leftarrow I$, $S\leftarrow S_{\textup{y}}$, $M\leftarrow\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{pos}}^{-1/2}$ to find
@@
where $w_{i}$ is the eigenvector corresponding to the eigenvalue $\lambda_{i}$ given by Proposition 3.4 and $\varphi_{i}$ is the right singular vector corresponding to $\lambda_{i}$ in (20). By [18, Theorem 2.8], the Moore–Penrose inverse of $\sum_{i=1}^{n}\sqrt{\frac{-\lambda_{i}}{1+\lambda_{i}}}w_{i}\otimes\varphi_{i}$ is given by $\sum_{i=1}^{n}\sqrt{\frac{1+\lambda_{i}}{-\lambda_{i}}}\varphi_{i}\otimes w_{i}$. Furthermore, the Moore–Penrose inverse of a composition of bounded operators is the composition in reverse order of the Moore–Penrose inverses of these operators, e.g. [23, eq. (3.23)]. Since $\mathcal{C}_{\textup{pr}}^{-1/2}$ and $I+\sum_{i=1}^{n}\frac{-\lambda_{i}}{1+\lambda_{i}}\varphi_{i}\otimes\varphi_{i}$ are boundedly invertible by Lemma A.7, it thus holds that the bounded operator
@@
has Moore–Penrose inverse equal to $S$. Because [2, Theorem 9.2(f)] implies that $(\mathfrak{S}^{\dagger})^{\dagger}=\mathfrak{S}$ for any bounded operator $\mathfrak{S}$, the operator in the display above is equal to $S^{\dagger}$. Furthermore, by [18, eq. (2.12)], $P_{\ker{S}^{\perp}}=S^{\dagger}S$, showing that $P_{\ker{S}^{\perp}}=\sum_{i=1}^{n}\varphi_{i}\otimes\varphi_{i}.$ Next, we compute for the given choice of $M$,
@@ B.3 Proofs of Section 7
Recall that H{H} defined in (2) is the Hessian of the negative log-likelihood of (1). Analogously, let $\widetilde{H}$ denote the Hessian of the negative log-likelihood of (29) with the given $P$. That is, upon replacement of $G$ with $GP$ in (2), we obtain $\widetilde{H}$. Hence, orthonormality of $(\varphi_{i})_{i}$ implies
@@ Appendix C Examples
Let us now assume that $\kappa$ is bounded and symmetric, and satisfies $\int\kappa(s,t)h(s)h(t)\geq 0$ for all $h\in\mathcal{H}$. Hence, $T_{\kappa}$ is self-adjoint and nonnegative. Then by Mercer’s theorem, [24, Theorem 3.a.1], we have $\kappa(s,t)=\sum_{i=1}^{\infty}b_{i}f_{i}(s)f_{i}(t)$, where the series converges absolutely and uniformly for almost every $(t,s)$. Here, $(b_{i})_{i}$ is a nonnegative sequence converging to zero and $(f_{i})_{i}$ is an ONB of $\mathcal{H}$ consisting of bounded functions. Furthermore, we may write $T_{\kappa}=\sum_{i}b_{i}f_{i}\otimes f_{i}$. For simplicity, we assume that the eigenvectors $(e_{i})_{i}$ of the prior covariance and the eigenfunctions $(f_{i})_{i}$ of the kernel are the same. One can verify that, with $a_{k,j}\coloneqq\langle f_{k},1_{[t_{j},t_{j+1}]}\gamma\rangle$, we have $\langle T_{\kappa}h,1_{[t_{i},t_{i+1}]}\gamma\rangle=\sum_{j}b_{j}a_{j,i}\langle f_{j},h\rangle$ and $\int\kappa(t,\cdot)1_{[t_{i},t_{i+1}]}(t)\gamma(t)\operatorname{d}\!{t}=\sum_{k}b_{k}a_{k,i}f_{k}$. Thus, $\mathcal{C}_{\textup{pr}}^{1/2}G^{*}\mathcal{C}_{\textup{obs}}^{-1/2}z=\sum_{i=1}^{n}\sum_{j}z_{i}b_{j}c_{j}a_{j,i}f_{j}$ for $z\in\mathbb{R}^{n}$. Furthermore, $G^{*}G=\sum_{i=1}^{n}\sum_{j,k}b_{j}b_{k}a_{j,i}a_{k,i}f_{k}\otimes f_{j}$ and hence the prior-preconditioned Hessian now takes the form
@@
$\partial_{t}u-\partial_{xx}u$ $=0,$ $\text{in }(0,1)\times(0,T),$
$u(\cdot,0)$ $=x^{\dagger},\quad$ $\text{on }(0,1),$
$u(0,\cdot)=u(1,\cdot)$ $=0,$ $\text{on }(0,T],$
where the true initial condition $x^{\dagger}$ is unknown and where we impose a homogenous Dirichlet spatial boundary condition. We assume that the data consists of a noisy observation of $u$ at the observation coordinates $(x_{i},t_{i})_{i=1}^{n}\subset(0,1)\times(0,T]$, where we assume i.i.d. standard Gaussian noise. The aim is to reconstruct the initial condition $x^{\dagger}$ from the data $y$. This problem is similar to [40, Example 3.5] and [20, Section 4.2], but in this example we do not observe the temperature field over the entire spatial domain at finitely many times. Instead, we observe the temperature only at finitely many space-time points $(x_{i},t_{i})_{i=1}^{n}$. Furthermore, [20, Section 4.2] considers periodic boundary conditions instead of Dirichlet boundary conditions. We take the Bayesian perspective by considering $x^{\dagger}$ as an $\mathcal{H}$-valued random variable $X$ with centered Gaussian distribution $\mu_{\textup{pr}}$. Below, we choose an explicit form of the prior covariance $\mathcal{C}_{\textup{pr}}$ as a negative power of the Laplacian.
To write this problem in the formulation of Section 2, we define $\mathcal{H}\coloneqq L^{2}((0,1))$. Let us denote by $H^{1}((0,1))$ the Sobolev space of square-integrable functions $h$ on $(0,1)$ that have a square-integrable weak derivative $\partial_{x}h$, which is a Hilbert space with the inner product $\langle h_{1},h_{2}\rangle_{1}\coloneqq\langle h_{1},h_{2}\rangle+\langle\partial_{x}h_{1},\partial_{x}h_{2}\rangle$, $h_{1},h_{2}\in H^{1}((0,1))$. By [19, Theorem 5.6.5], we have the continuous embedding $H^{1}((0,1))\subset C([0,1])$, where $C([0,1])$ denotes the space of continuous functions on $[0,1]$ with the supremum norm. Hence, for any $h\in H^{1}((0,1))$ and $x\in[0,1]$, we have $\lvert h(x)\rvert\leq\lVert h\rVert_{C([0,1])}\leq c\lVert h\rVert_{1}$ for some $c>0$, so that pointwise evaluation is well-defined, linear and continuous on $H^{1}((0,1))$. Thus, $H^{1}((0,1))$ is a reproducing kernel Hilbert space. We denote the Riesz representatives of the pointwise evaluation functionals, or ‘features’, by $\{\phi(x)\in H^{1}((0,1)),\ x\in[0,1]\}$. Hence, $h(x)=\langle h,\phi(x)\rangle_{1}$ for all $x\in[0,1]$ and $h\in H^{1}((0,1))$. For our choice of spatial domain $(0,1)$, we have the following explicit form for the features, by [41, Corollary 2]:
@@
We use certain properties of $\Delta\coloneqq\partial_{xx}$, the one-dimensional Laplacian. We describe these briefly, and refer to [25, Section 5.3] for a comprehensive treatment of these properties and their relation to the heat equation. By [5, Theorem 8.22], we can write $\Delta h=-\sum_{i}a_{i}\langle h,e_{i}\rangle e_{i}$ for $h\in\operatorname{dom}{\Delta}=\{h\in L_{2}((0,1)):\ \sum_{i}a_{i}^{2}\langle h,e_{i}\rangle^{2}<\infty\}$, where $\lim_{i}a_{i}=\infty$ and $(e_{i})_{i}$ is an ONB on $\mathcal{H}$. In fact, by the example on [5, p. 232], we have $a_{i}=i^{2}\pi^{2}$ and $e_{i}(x)=\sqrt{2}\sin(i\pi x)$ for our choices of spatial domain $(0,1)$ and boundary conditions. Now, one can define the self-adjoint operator $\exp(t\Delta)\in\mathcal{B}_{0}(\mathcal{H})$ by $\exp(t\Delta)=\sum_{i}\exp(-ta_{i})e_{i}\otimes e_{i}$. It holds that $\ker{\exp(t\Delta)}=\{0\}$ and $\operatorname{ran}{\exp(t\Delta)=\mathcal{H}}$. The diagonalisation of the Laplacian is compatible with $H^{1}_{0}((0,1))$ in the sense that $H^{1}_{0}((0,1))=\{h\in\mathcal{H}:\ \sum_{i}a_{i}\langle h,e_{i}\rangle^{2}<\infty\}$ and $\langle h,k\rangle_{1}=\sum_{i}(1+a_{i})\langle h,e_{i}\rangle\langle e_{i},k\rangle$ for $h,k\in H^{1}_{0}((0,1))$. Since for any $t\in(0,T]$ and $h\in\mathcal{H}$, we have $\langle\exp(t\Delta)h,e_{i}\rangle=\exp(-a_{i}t)\langle h,e_{i}\rangle$, it follows that $\sum_{i}a_{i}\langle\exp(t\Delta)h,e_{i}\rangle^{2}\leq C(t)\sum_{i}\langle h,e_{i}\rangle^{2}$ for some $C(t)>0$, so that $\exp(t\Delta)h\in\mathcal{H}^{1}_{0}((0,1))$. Therefore, the map $h\mapsto\exp(\Delta t)h$, $\mathcal{H}\rightarrow H^{1}_{0}((0,1))$ is linear and continuous for $t\in(0,T]$. Furthermore, $\exp(\Delta t)\in\mathcal{B}(H^{1}_{0}((0,1)))$ for each $t\in[0,T]$, and $\exp(\Delta t)$ is a self-adjoint element of $\mathcal{B}(H^{1}_{0}((0,1)))$, because